If P(A)=0.6, P(B)=0.3 and the events A and B are independent, then P(A∪ B) is
Single correct option
If P(A)=0.6, P(B)=0.3 and the events A and B are independent, then P(A∪B) is
A.0.9
B.0.18
C.0.42
D.0.72
Show answer and solution
Answer:(D) 0.72
For independent events P(A∩B)=P(A)P(B)=0.18. Then P(A∪B)=P(A)+P(B)−P(A∩B)=0.6+0.3−0.18=0.72. The value 0.9 simply adds the probabilities (valid only for mutually exclusive events), 0.18 is the intersection, and 0.42 results from subtracting the product twice. Equivalently 1−0.4×0.7=0.72.